OpenStudy (anonymous):

Tyrone almost forgot that he promised to pick up a prescription for his mother at the drug store across town. His mother is due home in 45 minutes. If Tyrone is able to make the 12.5 mile trip to the store and back again before his mother gets home, can he obey the posted speed limit of 35 mph?

OpenStudy (anonymous):

Total distance that needs to be traveled for the trip and back = 12.5 +12.5 miles = 25 miles. So,\[Distance = 25 miles\]He needs to make the trip in 45 minutes\[Time=45 minutes\]\[Time=\frac{45}{60}Hours = 0.75 Hours\]Use \[Speed = \frac{Distance}{Time}\]and find the speed and check if it is less than the posted speed limit of 35 miles/Hour.